# EVQAScore: Efficient Video Question Answering Data Evaluation

### v2

@@ 3.4 Video QA Evaluation
Uniform (Interval=30) 464 25.70 33.41
Uniform (Interval=20) 927 25.72 33.43
Uniform (Interval=10) 1857 25.73 33.44
None 18504 25.67 33.37
Table 2: Performance Comparison on VATEX-EVAL Dataset (No Reference). The table presents the Kendall $\tau_{b}$ and Spearman $\rho$ values, comparing Emscore, PAC-Score, and our proposed Emscore+. Emscore+ achieves the best performance, as indicated by the improvement values in parentheses.
@@
EVQAScore(PAC-S++) 32.8 42.3
(+4.7) (+5.9)
Table 3: Performance of VideoLLaVA and VideoLLaMA2 on ActivityNet, MSRVTT, MSVD, and TGIF shows that with only 12.5% of the data, EVQAScore outperforms the baseline using 100% of the data. Additionally, our approach surpasses PAC-S on both models across all evaluation datasets.
@@ 4 Experiments
In this section, we employ our Keyprompt method on both short and long benchmarks. We aim to address the following questions to verify the effectiveness of our method:
Q1: Can our EVQAScore achieve efficient computation without compromising evaluation performance?
Q2: Does our EVQAScore outperform all baseline methods in video caption data evaluation?
Q3: Does our EVQAScore outperform all baseline methods in selecting Video QA pairs?
Q4: Does our EVQAScore demonstrate superior performance compared to baseline methods when training VideoLLMs?
Q5: Can we visualize the advantages of our EVQAScore relative to previous video caption evaluation methods?
@@ 4.1 Experimental Setting
For the QA evaluation, we selected 200K samples from two datasets: the VideoInstruct100K [27] dataset and the VCG+ [28] 112K dataset. We then augmented these 200K samples with an additional 200K low-quality (noisy) data entries created by substituting each answer with a single keyword extracted from the answer sentence.
@@
To assess the alignment of our metric with human judgment, we utilized the VATEX-EVAL dataset to evaluate the correlation between our scores and human ratings in video captioning tasks. The VATEX-EVAL dataset is specifically designed to measure the correlation between automatic evaluation metrics and human judgment. It consists of 3,000 validation videos from the VATEX dataset, with a total of 18,000 candidate captions, six for each video.
@@
For LLaMA3.1-8B-Instruct, we applied the official hyperparameters from the repository. For the CLIP model, we used the clip-ViT-B-32 variant for keyframe selection. All experiments were conducted on a machine equipped with 8 NVIDIA H800 GPUs, a 192-core CPU, and 960GB of memory.
Table 4: Comparison of Noisy data filtered by PAC-S and our EVQAScore. We select much less noisy data.
Extract 25% data
Method noisy videochatgpt videochatgpt+ total
PAC-S 27928 40012 33764 101704
EVQAScore 1101 53341 47262 101704
Extract 12.5% data
PAC-S 10832 21095 18925 50852
EVQAScore 450 25642 25165 50852
4.2 Computational Efficient EVQAScore
To address Q1, we conduct uniform frame extraction in EVQAScore. We select the sampling interval to be 10, 20 and 30 frames. The results is presented in Table 1, they illustrate that using uniform keyframe extraction substantially reduces the time required for managing video data, with significantly shorter processing times compared to not extracting keyframes. For instance, applying a uniform interval of 30 frames results in a processing time of 464 seconds, compared to 18,504 seconds for no keyframe extraction, representing a drastic time reduction. Importantly, this improvement in efficiency comes with nearly no impact on performance, as evidenced by the Kendall and Spearman correlation scores, which remain virtually unchanged across different intervals. This demonstrates that uniform frame selection is a highly efficient approach, achieving considerable time savings without compromising the evaluation quality.
4.3 EVQAScore for Video Caption Evaluation
To address Q2, we conducted video caption evaluation experiments on the VATEX-EVAL dataset. The results are presented in Table 2. Our findings demonstrate that EVQAScore consistently outperforms both EMScore and PAC-S in correlating with human judgment. Specifically, EVQAScore achieves a Kendall $\tau_{b}$ score of 25.6 and a Spearman $\rho$ score of 33.3 in the No Reference setting using the PAC-S CLIP model, representing an improvement of 0.5 points in Kendall and 0.7 points in Spearman over PAC-Score. Additionally, further experiments using PAC-S++ checkpoints resulted in SOTA performance, with a Kendall $\tau_{b}$ score of 32.8 and a Spearman $\rho$ score of 42.3, marking improvements of 4.7 points and 5.9 points over the PAC-S++ baseline. These results indicate that our proposed method better aligns with human perception of video-caption consistency in the No Reference evaluation setting.
This strong performance highlights EVQAScore’s capacity to enhance existing scoring systems. It demonstrates that EVQAScore excels not only in evaluating Video QA data but also in video caption data evaluation. This further underscores the potential of EVQAScore for high-fidelity video caption data assessment. As our method utilizes only keywords, previous caption-based methods may suffer from the token limitation (77 tokens) of CLIP, making our approach more suitable for evaluating very long videos and their comprehensive caption. This strong performance highlights EVQAScore’s capacity to amplify existing scoring systems, further emphasizing its potential for high-fidelity video QA data evaluation.
4.4 Video QA Data Filtering
Table 5: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench-Diverse.
@@ 4.4 Video QA Data Filtering
EVQAScore 25% 25% 1.75 2.27 2.43 0.86 2.07 1.88 1.05 1.94 3.04
EVQAScore 12.5% 12.5% 1.76 2.39 2.47 0.83 1.97 1.88 1.00 1.98 3.25
Table 6: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench.
Method Models Data Proportion CI DO CU TU CO Avg.
Baseline VideoLLaMA2 100% 0.60 0.29 0.88 0.21 1.05 0.61
PAC-S 25% 2.02 1.70 2.35 1.59 2.48 2.03
PAC-S 12.5% 2.39 2.15 2.75 1.73 2.75 2.35
EVQAScore 25% 3.01 2.98 3.38 2.33 3.05 2.95
EVQAScore 12.5% 3.02 3.14 3.39 2.23 3.09 2.97
Baseline VideoLLaVA 100% 0.59 0.30 0.87 0.20 0.96 0.58
PAC-S 25% 2.18 1.96 2.53 1.60 2.54 2.16
PAC-S 12.5% 2.24 2.08 2.63 1.82 2.60 2.27
EVQAScore 25% 2.77 2.82 3.15 2.16 3.04 2.79
EVQAScore 12.5% 2.68 2.93 3.09 2.15 2.71 2.71
Table 7: Performance of VideoLLaVA and VideoLLaMA2 on MVBench using filtered data by PAC-S and EVQAScore.
Method Model Data Proportion Avg.
Baseline VideoLLaMA2 100% 34.09
PAC-S 25% 25% 31.43
PAC-S 12.5% 12.5% 34.24
EVQAScore 25% 25% 34.89
EVQAScore 12.5% 12.5% 37.60
Baseline VideoLLaVA 100% 30.68
PAC-S 25% 25% 37.10
PAC-S 12.5% 12.5% 37.53
EVQAScore 25% 25% 38.01
EVQAScore 12.5% 12.5% 38.96
To address Q3, we perform data filtering using our EVQAScore metric alongside the previous state-of-the-art (SOTA) data evaluation baseline, PAC-S. The results presented in Table 4 illustrate the effectiveness of the EVQAScore metric in filtering Video QA data. For both the top 25% and 12.5% data extractions, EVQAScore significantly reduces noisy data while retaining a higher proportion of VideoChatGPT and VideoChatGPT+ data compared to the PAC-S method. Specifically, EVQAScore retains only 1,101 and 450 noisy data points for the 25% and 12.5% extractions, respectively, while preserving a greater volume of high-quality data. This indicates that EVQAScore is more selective, thereby enhancing data quality and processing efficiency for Video QA tasks.
By reducing the amount of low-quality data, we are able to train more effective VideoLLMs. Further experiments presented in Section 4.5 demonstrate the impact of EVQAScore on VideoLLM performance.
4.5 Performance of Video-QA Evaluation
To address Q4, we train two models VideoLLaVA and VideoLLaMA2 on our filtered datasets. Then we conduct comprehensive evaluation on several benchmarks, i.e. Activitynet, MSVD, MSRVTT, TGIF, MVBench, VCGBench and VCGBench Diverse. Our EVQAScore achieves SOTA performance across diverse models and benchmarks.
@@ 4.5 Performance of Video-QA Evaluation
Figure 3: A case study on evaluating very long QA pairs demonstrates that EVQAScore can effectively assess such cases, whereas CLIPScore, EMScore, and PAC-S struggle to do so.
We selected two different models: VideoLLaVA and VideoLLaMA2. These models have distinct architectures and use different encoders and foundation models. Specifically, VideoLLaVA employs LanguageBind as its encoder and Vicuna 1.5 as the foundation model, while VideoLLaMA2 utilizes CLIP as its encoder and LLaMA2 as the foundation model. Across the four benchmarks—ActivityNet, MSVD, MSRVTT, and TGIF presented in Table 3, VCGBench in Table 6, VCGBench-Diverse in Table 5, and MVBench in Table 7—our filtered data outperformed the previous SOTA baseline, PAC-S, for both model architectures. This demonstrates the robustness of our EVQAScore.
In this section, we demonstrate that our method effectively filters high-quality data to enhance VideoLLM performance. We showed that EVQAScore can improve the performance of models with different architectures across all benchmarks, underscoring the wide application potential of our EVQAScore.
4.6 Case Study: Evaluating Video Captions
To address Q5, we present case study, we demonstrate the effectiveness of EVQAScore using two cases.
EVQAScore can correctly evaluate comprehensive answers
We summarize the first case in Figure 1. Given the question: ”What is the man doing in the video?” the first answer contains only one word and is of low quality, while the second answer is longer and of higher quality. Compared with CLIPScore (using the pooling feature of video), EMScore, PAC-S, and EVQAScore, the video QA pair indicates that our method is the only approach capable of recognizing the second answer as being of higher quality. This demonstrates the robustness of our method in evaluating Video QA data.
EVQAScore can evaluate longer QAs
We summarize the second case in Figure 3. From the figure, we observe a man playing bagpipes. When asked to describe the content of the video, the first answer, ”pipes,” is too short and incorrect. In contrast, the second answer is comprehensive and considerably longer. However, due to the absence of keyword extraction, methods such as CLIPScore (using the pooling feature of video), EMScore, and PAC-S suffer from the limited token count (77 tokens) imposed by CLIP, requiring them to truncate the caption, thereby preventing a full understanding of the QA. Consequently, CLIPScore, EMScore, and PAC-S fail to demonstrate that Answer 2 outperforms Answer 1.
In this section, we use two real cases to illustrate the effectiveness of EVQAScore. In recent years, understanding longer videos has become increasingly essential, and evaluating the quality of long videos is a critical task. When applied to longer videos—often spanning hours and involving thousands of words in QA or captions—our method demonstrates a higher potential for success.
@@ 5 Conclusion
In this work, we address the critical challenge of evaluating video QA data quality by introducing EVQAScore, a reference-free evaluation method that leverages keyword extraction to assess both video caption and video QA data quality. By incorporating frame sampling and rescaling techniques, our method effectively handles the evaluation of extremely long videos, enhancing efficiency and robustness. EVQAScore achieves state-of-the-art performance on the VATEX-EVAL benchmark, with scores of 32.8 for Kendall correlation and 42.3 for Spearman correlation in video caption evaluation. Moreover, by applying EVQAScore for data selection, we demonstrate its effectiveness in selecting much less noisy data. Moreover, we achieved SOTA results, surpassing the previous SOTA method PAC-S and 100% data with only 12.5% of the original data volume, highlighting the potential for substantial improvements in data efficiency and overall video understanding.
Supplementary Material
6 Keywords Extraction Prompt
Figure 4: We provide the prompt used for keywords extraction

### v3

@@ 3.4 EVQAScore for VideoQA Data Quality Evaluation
EVQAScore (Interval=60) 2283.9 28.9 37.4
EVQAScore (Interval=50) 2582.6 29.0 37.5
EVQAScore (Interval=40) 3177.0 29.0 37.6
EVQAScore (Interval=30) 3986.0 29.1 37.6
EVQAScore (Interval=20) 5769.3 29.1 37.6
EVQAScore (Interval=10) 9774.6 29.1 37.6
EMScore 887.4 23.2 30.3
PAC-S 927.9 25.1 32.6
PAC-S++ 998.6 28.1 36.6
Table 2: Performance Comparison on VATEX-EVAL Dataset (No Reference). The table presents the Kendall $\tau_{b}$ and Spearman $\rho$ values, comparing Emscore, PAC-Score, and our proposed EVQAScore. EVQAScore achieves the best performance.
@@
EVQAScore(Alpha-CLIP) 26.6 34.5
EVQAScore(PAC-S) 26.4 34.3
EVQAScore(PAC-S++) 28.9 37.4
Table 3: Performance of VideoLLaVA and VideoLLaMA2 on ActivityNet, MSRVTT, MSVD, and TGIF. The results demonstrate that, with only 12.5% of the data filtered by EVQAScore, the VideoLLMs outperform the baseline using 100% of the data. Furthermore, our approach outperforms PAC-S in data filtering for both models across all evaluation datasets.
@@ 4 Experiments
In this section, we apply our EVQAScore method to both video-caption and videoQA datasets. The goal is to address the following questions to assess the effectiveness of our method:
Q1: Does our EVQAScore outperform all baseline methods in video-caption data evaluation?
Q2: Does our EVQAScore outperform all baseline methods in selecting VideoQA pairs?
Q3: Does our EVQAScore demonstrate superior performance compared to baseline methods when training VideoLLMs?
Q4: Can we visualize the advantages of our EVQAScore relative to previous video-caption evaluation methods?
@@ 4.1 Experimental Setting
For Video-Caption data evaluation, to assess the alignment of our metric with human judgment, we utilized the VATEX-EVAL dataset to evaluate the correlation between our scores and human ratings in video-captioning tasks. The VATEX-EVAL dataset is specifically designed to measure the correlation between automatic evaluation metrics and human judgment. It consists of 3,000 validation videos from the VATEX dataset, with a total of 18,000 candidate captions, six for each video.
For the VideoQA evaluation, we selected 200K samples from two datasets: the VideoInstruct100K [23] dataset and the VCG+ [24] 112K dataset. We then augmented these 200K samples with an additional 200K low-quality (noisy) data entries created by substituting each answer with keywords extracted from the answer sentence.
@@
For the VATEX-EVAL dataset, we utilize the Kendall $\tau$ and Spearman correlation coefficients to assess the relationship between our rating scores and human ratings.
@@
Table 4: We compare the noisy data filtered by PAC-S and our EVQAScore. EVQAScore selects significantly less noisy data compared to the previous state-of-the-art method, PAC-S.
Method noisy videochatgpt videochatgpt+ total
Extract 25% data
PAC-S 27928 40012 33764 101704
EVQAScore 456 53251 47997 101704
Extract 12.5% data
PAC-S 10832 21095 18925 50852
EVQAScore 15 24640 26197 50852
@@
For LLaMA3.1-8B-Instruct, we applied the official hyperparameters from the repository. For the CLIP model, we used the same CLIP score as EMScore [29], PAC-S, and PAC-S++ [28]. All experiments were conducted on a machine equipped with 8 NVIDIA H20 GPUs, a 192-core CPU, and 960GB of memory. For YOLO, we used the yolo11x-seg checkpoint and the default settings for key object extraction.
4.2 EVQAScore for Video-Caption Evaluation
To address Q1, we conducted video-caption evaluation experiments on the VATEX-EVAL dataset. The results are presented in Table 2. Our findings demonstrate that EVQAScore consistently outperforms both EMScore and PAC-S in correlating with human judgment. Specifically, EVQAScore achieves a Kendall $\tau_{b}$ score of 26.4 and a Spearman $\rho$ score of 34.3 in the No Reference setting using the PAC-S CLIP model, representing an improvement of 1.3 points in Kendall and 1.7 points in Spearman over PAC-Score. Additionally, further experiments using PAC-S++ checkpoints resulted in SOTA performance, with a Kendall $\tau_{b}$ score of 28.9 and a Spearman $\rho$ score of 37.4, marking improvements of 0.8 points and 1.0 points over the PAC-S++ baseline. These results indicate that our proposed method better aligns with human perception of video-caption consistency in the No Reference evaluation setting.
Table 5: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench-Diverse. Our EVQAScore filtered data achieved SOTA performance training VideoLLMs.
@@ 4.2 EVQAScore for Video-Caption Evaluation
EVQAScore 25% 25% 1.79 2.34 2.48 0.90 2.02 1.91 1.06 2.00 3.14
EVQAScore 12.5% 12.5% 1.80 2.53 2.55 0.83 2.05 1.95 1.03 2.08 3.39
Table 6: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench and MVBench. Our EVQAScore filtered data achieved SOTA performance training VideoLLMs.
Method Model Data Proportion VCGBench MVBench
CI DO CU TU CO Avg. Avg.
Baseline VideoLLaMA2 100% 0.60 0.29 0.88 0.21 1.05 0.61 34.09
PAC-S 25% 2.02 1.70 2.35 1.59 2.48 2.03 31.43
PAC-S 12.5% 2.39 2.15 2.75 1.73 2.75 2.35 34.24
EVQAScore 25% 3.06 3.09 3.45 2.30 3.11 3.00 33.34
EVQAScore 12.5% 3.03 3.31 3.41 2.36 3.17 3.06 36.65
Baseline VideoLLaVA 100% 0.59 0.30 0.87 0.20 0.96 0.58 30.68
PAC-S 25% 2.18 1.96 2.53 1.60 2.54 2.16 37.10
PAC-S 12.5% 2.24 2.08 2.63 1.82 2.60 2.27 37.53
EVQAScore 25% 2.79 2.88 3.19 2.22 2.92 2.80 39.50
EVQAScore 12.5% 2.66 3.03 3.08 2.05 2.60 2.68 38.69
This strong performance highlights EVQAScore’s capacity to enhance existing scoring systems. It demonstrates that EVQAScore excels in evaluating video-caption data evaluation.
4.3 EVQAScore for VideoQA Data Filtering
To address Q2, we perform data filtering using the EVQAScore metric in conjunction with the previous SOTA data evaluation baseline, PAC-S. The results presented in Table 4 demonstrate the effectiveness of the EVQAScore metric in filtering VideoQA data. For both the top 25% and 12.5% data extractions, EVQAScore significantly reduces noisy data while retaining a larger proportion of high-quality VideoChatGPT and VideoChatGPT+ data compared to the PAC-S method. Specifically, EVQAScore retains only 456 and 15 noisy data points for the 25% and 12.5% extractions, respectively, while preserving a higher volume of high-quality data. This indicates that EVQAScore is more selective, thereby enhancing both data quality and processing efficiency for VideoQA tasks.
By reducing the amount of low-quality data, we enable the more efficient training of higher-performing VideoLLMs. Further experiments, presented in Section 4.4, demonstrate the impact of EVQAScore on VideoLLM performance.
4.4 EVQAScore can Filter High-Quality VideoQA for VideoLLMs Training
To address Q3, we train two models with different structures VideoLLaVA and VideoLLaMA2 on our filtered datasets. Then we conduct a comprehensive evaluation on several benchmarks, i.e. Activitynet, MSVD, MSRVTT, TGIF, MVBench, VCGBench and VCGBench Diverse. Our EVQAScore achieves SOTA performance across diverse models and benchmarks.
@@ 4.4 EVQAScore can Filter High-Quality VideoQA for VideoLLMs Training
Figure 4: In this case study, Answer1 is more accurate, while Answer2 introduces hallucination. Our keyword extraction method successfully identifies hallucinations, making EVQAScore the only metric capable of determining that Answer1 is superior to Answer2.
We selected two different models: VideoLLaVA and VideoLLaMA2. These models have distinct architectures and use different encoders and foundation models. Specifically, VideoLLaVA employs LanguageBind as its encoder and Vicuna 1.5 as the foundation model, while VideoLLaMA2 utilizes CLIP as its encoder and LLaMA2 as the foundation model. Across the four benchmarks—ActivityNet, MSVD, MSRVTT, and TGIF presented in Table 3, VCGBench and MVBench in Table 6, VCGBench-Diverse in Table 5—our filtered data outperformed the previous SOTA baseline, PAC-S, for both model architectures. This demonstrates the robustness of our EVQAScore.
Overall, we demonstrate that our EVQAScore effectively filters high-quality data to enhance VideoLLM performance. We show that EVQAScore not only evaluates video-caption data quality effectively, as detailed in section 4.2, but also improves the quality of VideoQA data and, consequently, the performance of VideoLLMs across various architectures and benchmarks. This underscores the broad applicability of our EVQAScore.
4.5 Case Study: Evaluating VideoQA Data
To address Q5, we present a Case Study to demonstrate the effectiveness of EVQAScore. As shown in Figure 4, the user queries about the action depicted in the video. Answer1 provides a more accurate response, while Answer2 introduces hallucinations. Notably, our keyword extraction method successfully identifies the hallucinated part of Answer2, helping to distinguish it from the correct answer. This highlights the ability of EVQAScore to not only evaluate the relevance and accuracy of the answer but also to detect and penalize hallucinations. Therefore, EVQAScore correctly assigns a higher score to Answer1, reaffirming its superiority over Answer2.
@@ 5 Conclusion
In this work, we present EVQAScore, a reference-free evaluation method designed to assess the quality of both video-caption and VideoQA data for training VideoLLMs. By integrating patch retrieval, keyword extraction, and frame sampling, EVQAScore provides a fine-grained, efficient approach to video data quality evaluation. Our method surpasses existing techniques, achieving SOTA performance on the VATEX-EVAL benchmark, with significant improvements in Kendall and Spearman correlation scores. Additionally, EVQAScore enables effective data selection, reducing the data volume by over 87.5% while maintaining high performance. The promising results demonstrate the potential of EVQAScore in enhancing the evaluation and selection of high-quality video data, contributing to the development of more efficient and accurate VideoLLMs. Our future work will explore further optimizations and broader applications of this method in other video-based tasks.